OpenStudy (anonymous):

choose the point on the terminal side of theta=2pi/3

OpenStudy (luigi0210):

Do you know your unit circle?

OpenStudy (luigi0210):

You have to move the it to \[\LARGE \theta=\frac{2 \pi}{3} \] So:  Now find the points that exist there. Which would be \[\LARGE (-\frac{1}{2},~\frac{\sqrt{3}}{2})\] And yea, there ya go.

OpenStudy (anonymous):

oh, thank you so much, i was overcomplicating it!

OpenStudy (luigi0210):

You're welcome, and \[\LARGE \color{blue}{WELCOME~~TO~~OPENSTUDY} \]

OpenStudy (luigi0210):

So if you have any other questions(relating to the site itself), feel free to PM me or any other person with A next to their name/enclosed in purple. I hope you enjoy your time on here. :) And just a head's up for future reference, you can give away medals to people who have help you(if you feel they deserve it) by pressing the "Best Response" button next to their name.